You're driving along at 25 m/s with your aunt's va
Answer #1

V^2 = U^2 + 2*a*S

0 = 25*25 - 2*a*55

=> a = 5.68 m/s^2

Yes we can stop the truck if retardation of truck is more than 5.68 m/s^2